The following data represents the amount in kg of fertilizer applied to equal size of plots and the yields in kg of maize.

PLOT

AMOUNT OF FERTILIZER(X)

YIELDS(Y)

A

2

7

B

1

3

C

3

8

D

4

10

 

(a)   Calculate and interpret the correlation coefficient for X and Y               6 marks

 

(b)  Calculate and interpret the least square regression of Y on X                 8 marks

 


1
Expert's answer
2021-12-27T10:53:13-0500

a)

correlation coefficient:

"r =\\frac{\t\u03a3(x_i - x\u0304)(y_i - \u0233)}{\\sqrt{(\u03a3(x_i - x\u0304)^2\u03a3(y_i - \u0233)^2 )}}=0.9648"

there is a significant large positive relationship between X and Y


b)

regression line:

"y=b_1x+b_0"

where

"b_1=\frac{Σ(x_i-x̄)(y_i-ȳ)}{Σ(x_i-x̄)^2}=2.2"


"b_0 = \u0233 - b_1x\u0304=1.5"


slope of line b1 is rate of change of yields (Y) respect to amount of fertilizer (X)

b0 is value of yields for amount of fertilizer = 0